OpenStudy (anonymous):

x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

find the LCM for each equation

OpenStudy (anonymous):

I have no idea what that means though. /: I don't understand these problems at all.

OpenStudy (matt101):

You're solving a system of equations with three equations and three unknowns. You probably know that in order to find the value of each variable, you can only have one variable in the equation. This means you have to express each variable in terms of the others and substitute...I'll show you what I mean: Take the first equation, x + 4y + z = –10. Isolating for x, you get x = -10 - 4y - z. This is x expressed in terms of y and z. Now you can take that and sub it into your second equation, 3x – 3y + 6z = –21. However, before you do, note that you can divide the entire equation by 3 just to make it a little easier to work with. Dividing by 3, you get x - y + 2z = -7. Now sub in what you found from the first equation for x: (-10 - 4y - z) - y + 2z = -7 After you simplify, you'll see that there are now only two variables in that equation. Isolate either one, then sub that and x into your third equation to solve for the remaining variable. Does that make sense?

OpenStudy (anonymous):

Yes! Thank you so much!(:

OpenStudy (matt101):

Np if you need me to clarify anything I'm around...